OpenStudy (anonymous):

Find the average value of the function on the given interval. F(theta)=sec^2(theta/2), [0, pi/2]

OpenStudy (anonymous):

\[\frac{\sec^2(\frac{\pi}{4})-\sec^2(0)}{\frac{\pi}{2}-0}\] is a start then you have to compute

OpenStudy (anonymous):

oh unless this means "mean value" in which case you have to integrate

OpenStudy (anonymous):

Nope it just says the average value.

OpenStudy (anonymous):

are you doing integral calculus?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

\(\sec(\frac{\pi}{4})=\sqrt{2}\) and \(\sec(0)=1\) so you would get \[\frac{2-1}{\frac{\pi}{2}}=\frac{2}{\pi}\]

OpenStudy (anonymous):

or if this is mean value you would get \[\frac{2}{\pi}\int_0^{\frac{\pi}{2}}\sec^2(\frac{x}{2})dx\]

OpenStudy (anonymous):

anti derivative of secant squared is tangent, so it should be easy to complete